OpenStudy (anonymous):

Simplify. 2 + 4 • (6 – 3) A. 18 B. 14 C. 23 D. 33

OpenStudy (igreen):

2 + 4 * (6 - 3) Parenthesis: 2 + 4 * 3 Multiplication: 2 + 12 Addition: 14

OpenStudy (anonymous):

...

OpenStudy (anonymous):

Just tell me the answer!

OpenStudy (anonymous):

First find your solution to what ever is in parenthesis. in this case, 6-3 which equals 3.

OpenStudy (igreen):

That's it..14..@JillyCraft64

OpenStudy (anonymous):

Thats not how it works. I will tell you how to get the answer, I wont just tell you the answer

OpenStudy (anonymous):

You aren't supposed to do that. :/

OpenStudy (anonymous):

Well some people say A. B. C. or D. Now that seems Easier

OpenStudy (igreen):

I can.

OpenStudy (anonymous):

It really easy, @JillyCraft64 Will you just let me show you howto get the answer?

OpenStudy (igreen):

@JillyCraft64 Maybe it does, but you're not supposed to give out answers WITHOUT explanation.

OpenStudy (igreen):

And don't you think it's better to have an explanation? That way you know the person has answered your question correctly.

OpenStudy (anonymous):

WELL im Clueless

OpenStudy (anonymous):

Let me show you.

OpenStudy (anonymous):

i know theres a d word but i cant spell it and DONT show me

OpenStudy (igreen):

You have to follow PEMDAS. P is Parenthesis, you do this first if there is a parenthesis. You have to follow PEMDAS in the Parenthesis, too. E is Exponent, you do this next. M and D is Multiplication and Division, you do whichever one comes first from left to right. A and S is Addition and Multiplication, you do whichever one comes first from left to right.

OpenStudy (igreen):

@JillyCraft64 Can you solve \(12 \div 3 \times 4~?\)

OpenStudy (anonymous):

16?

OpenStudy (igreen):

Yes! You divide first, then you multiply.

OpenStudy (igreen):

How about this one? \(6 + (3^2 - 4)\)

OpenStudy (anonymous):

...

OpenStudy (igreen):

Do you know what exponents are? @JillyCraft64

OpenStudy (igreen):

Have you learned about them yet..?

OpenStudy (anonymous):



OpenStudy (anonymous):

I URGE YOU TO READ THIS This is easier to understand. But if you look at it you will realize its supppeeerrr easy!!! I thought it was hard when I was learning how to do this (7th grade) but trust me, once you understand how to do this, its really easy. :) Tell me if you need anymore help! You should use the order of operations, pmdas (parenthesis, multiplication, division, addition, and last subtraction. 2+4*(6-3) IN the order of operations parenthesis are first...so 6-3= 3 Now our equation will look like this- 2+4*3... since we figured out (6-3) we replace the question with an answerand move on. 2+4*3 Since multiplication is next, we have to know what to multiply. The numbers to the left and to the right of the multiplication sign are 4 and 3 so... 4*3=12 = 12 Again, we will replace the question (4*3) with an answer(12). Now the equation looks like this--- 2+12 Division is next, there aren't any division signs in this equation, so we skip this step. The Then theres addition, there is one addition sign so... 2+12= 14... and that's! Your answer is 14. Reply Using Drawing

OpenStudy (anonymous):

that last sentence wasn't supposed to be there